# Highly Undersampled MRI Reconstruction via a Single Posterior Sampling of Diffusion Models

### v1

@@ 6 Discussion and Conclusion
In this work, we demonstrated the feasibility of MRI reconstructions from highly undersampled k-space with only a single-step posterior sampling of a diffusion model (DM). Specifically, an SSDM-MRI framework was constructed by progressively distilling a conventional DM. Extensive comparative studies were carried out to compare the performances of SSDM-MRI with those of multiple existing algorithms, i.e., P-LORAKS [49], CSPC [40], ADMM-Net [17], U-net [16], and DM-based Score-MRI [26] on both public fastMRI images and a private QSM subject. In general, the proposed method exhibited the best results on most images with fewer errors, improved numerical metrics and image fine details, and more accurate susceptibility measurements in the substantia nigra and red nucleus.
@@
There have already been several DM-based methods proposed for MRI acceleration [24, 25, 26, 27, 28, 29, 30, 31], showing better reconstruction performances, especially when compared with traditional iterative and existing end-to-end deep learning (DL) based methods [12, 13, 14]. However, the computational costs of DM-based methods are usually much higher than previous DL methods, e.g., U-net-based frameworks, due to the need of a relatively large number of poster sampling steps, making the overall reconstruction time much longer and hindering their applications in clinical practice. To the best of our knowledge, SSDM-MRI is the first single-step DM-based method, whose reconstruction time is comparable to the end-to-end deep neural networks, as evident in Table 1.
@@
In this work, the proposed SSDM-MRI demonstrated promising generalization capability on the OOD knee data, despite the fact that it was trained on brain images only, demonstrating SSDM-MRI’s ability to learn the conditional prior underlaid in shared anatomical and statistical patterns of brain and knee images instead of simply memorizing the features. Although the adaptability of SSDM-MRI may not be as good as fully unsupervised methods (e.g., the slight performance decrease in Tabel 2 for knee images), it still demonstrates great robustness across vairous anatomies.
@@
In the QSM experiment, different methods were repetitively tested on 1024 (128 slice number 8 echo number) consecutive 2D images, covering the whole brain before QSM reconstructions from MRI phases. The proposed SSDM-MRI not only achieved on average the best results for all brain slices, but also preserved the best slice-to-slice coherence. It also resulted in the best QSM measurements, demonstrating its capability of preserving latent and weak susceptibility signal buried in the MRI phase images.
@@
SSDM-MRI’s performances were also evaluated against varying accelerating rates, and the results showed that no substantial artifacts were introduced. However, the image quality also gradually degraded with increasing AFs, and in the future, we will continue working on obtaining high-quality images with high AFs. In addition, the reconstruction stability also needs to be further optimized to obtain more reliable solutions.

### v2

@@ 6 Discussion and Conclusion
In this work, we demonstrated the feasibility of MRI reconstructions from highly undersampled k-space with only a single-step posterior sampling of a diffusion model (DM). Specifically, an SSDM-MRI framework was constructed by iteratively distilling a pretrained conditional DM using the proposed iterative selective distillation algorithm (paired with a paired shortcut reverse sampling strategy). Extensive comparative studies were carried out to compare the proposed SSDM-MRI with multiple existing algorithms, i.e., P-LORAKS [11], CS [10], MoDL [26], U-net [23], and DM-based Score-MRI [33] and CCDF [40] on both public fastMRI images and a private multi-echo GRE (QSM) subject. In general, the proposed method exhibited the best results with the minimum reconstruction errors, the best numerical metrics and image fine details, and the most accurate susceptibility measurements in five deep grey matter regions (i.e., Substantia Nigra, Red Nucleus, Globus Pallidus, Caudate, and Putamen).
@@
There have already been several DM-based methods proposed for MRI acceleration [34, 41, 33, 37, 42, 43, 38, 39], showing better reconstruction performances, especially when compared with traditional iterative and existing end-to-end deep learning (DL) based methods [17, 18, 19]. However, the computational costs of DM-based methods are usually much higher than previous DL methods, e.g., U-net-based frameworks, due to the need of a relatively large number of poster sampling steps, making the overall reconstruction time much longer and hindering their applications in clinical practice. To the best of our knowledge, there have not been any DM-based methods that can achieve high-quality MRI reconstructions in just one posterior sampling step, and SSDM-MRI should be the first one in the MRI community, and its reconstruction time is only comparable to end-to-end deep neural networks, as is evident in Table 1.
@@
The proposed SSDM-MRI demonstrated promising generalization capability on the OOD knee data, though it was trained on brain images only, demonstrating its ability to learn the conditional prior underlaid in shared anatomical and statistical patterns of brain and knee images instead of simply memorizing the features.
It is found in the ablation experiment (Fig. 10) that the performance of the proposed SSDM-MRI decreases with increasing number of distillations, which means that there is a compromise between the model’s performance and the computational cost. Reducing the necessary number of distillations might be an effective way to handle the trade-off between SSDM-MRI’s performance and the computational costs. In this work, an iterative selective distillation combined with a shortcut reverse sampling framework was proposed to reduce the number of distillations and, in the meantime, to focus only on the most important parts of the reverse sampling, rather than the whole reverse path from the pure noise to clean images, which is the key design for the improved performance.
| PSNR | ||||||||||||
| SSIM |
A relatively simple and straightforward projection-based DC (PDC) method was adopted in the current study after the reconstruction of the model to further improve the results and reduce the variance in the DM results. More advanced DC methods can be incorporated to further improve the results. We have conducted an experiment to compare the PDC method with two more advanced DC methods, i.e., a two-step DC (TDC) [57], and a learning-based DC (LDC) [58] on retrospectively undersampled brain images, with the results reported in Fig. 11 (one brain slice undersampled at 12) and Table 4 (128 brain images undersampled at AF = 8). Overall, SSDM-MRI with TDC or LDC achieved visually similar results and slightly better quantitative metrics (PSNR and SSIM), compared with SSDM-MRI using PDC. However, these improvements are much smaller than the improvements from SSDM-MRI without DC to SSDM-MRI with PDC. For example, as shown in Table 4, changing PDC to LDC or TDC only results in an average 0.17/0.005 or 0.07/0.019 increase in PSRN/SSIM, respectively, compared to 1.01/0.033 increase from SSDM-MRI without data consistency to SSDM-MRI with the simplest PDC. In addition, PDC is more generic than LDC and TDC in some specific scenarios. For example, the LDC will need to be retrained to better suit different methods, while TDC cannot be directly used on the mGRE data reconstructions because it discards the very important phase component after the first DC step. Overall, the choice of DC methods can be very flexible if it is applied fairly to all comparative methods.
We also performed another experiment to compare multi-coil image reconstructions based on SVD and ESPIRiT coil combination in Fig. 12. It is obvious that the proposed SSDM-MRI achieved the best results compared to all other MRI reconstruction methods, regardless of which coil sensitivity estimation method is applied. In addition, it is demonstrated that the SVD-based method yields slightly higher PSNRs and SSIMs compared to the results obtained with the ESPIRiT method for all comparative methods. Therefore, the main conclusions will remain unchanged using either method.
@@
In the mGRE data-based experiment (i.e., the QSM acceleration experiment), different methods were repetitively tested on 1024 (128 slice number 8 echo number) consecutive 2D images, covering the whole brain, before QSM and R2* reconstructions from the reconstructed phases and magnitude images, respectively. The proposed SSDM-MRI not only achieved on average the best results for all brain slices, but also preserved the best slice-to-slice coherence. It resulted in the best QSM measurements, demonstrating its capability of preserving latent and weak susceptibility signals buried in the MRI phase images. In addition, SSDM-MRI also achieved the most accurate results in R2*-based comparison, which further demonstrates its applicability in multi-echo imaging contexts.
@@
SSDM-MRI’s performances were also evaluated against varying accelerating rates, and the results showed that no substantial artifacts were introduced. However, the image quality also gradually degraded with increasing AFs, and in the future, we will continue working on obtaining high-quality images with high AFs. In addition, the reconstruction stability also needs to be further optimized for more reliable solutions.